# Towards a Method for LLM-enabled Indoor Navigation (Arxiv) thanks: This work was founded by European Union - Next Generation EU, in the context of The National Recovery and Resilience Plan, Investment Partenariato Esteso PE8 ”Conseguenze e sfide dell’invecchiamento”, Project Age-IT, CUP: B83C22004880006

### v2

@@ III Experimental Set-up
IV Results and Discussion
In the following, we comment on the results for each map, as each case has specific characteristics and challenges, which allow us to point out the strengths and weaknesses of the tested LLM.
IV-A Bergamo
The image of the Bergamo airport is very well-colored and clearly represented. It includes various IDs placed at specific locations on the map and additional icons that highlight points of interest, such as staircases, access points, phone charging stations, or information desks. The layout is characterized by a single corridor with a few distracting elements in the middle, and all various shops or restaurants appear along the sides. However, there are many services and points of interest to recognize, and, often, the IDs are placed very close to each other or are very small. This complicates the correct parsing of image information by the LLM.
The detailed results for Bergamo Airport are summarized in Table I. Over half of the indications are incorrect (62.26%), while 37.74% are correct.
IV-B Bologna
Similar to Bergamo airport, this map features a structure defined by a large, single corridor with shops on the sides. It differs from the first one in its use of colors, the overall representation of the floor, and a less linear configuration, with a much more pronounced central curve.
Table II presents the detailed results for Bologna. Out of 66 indications, 27 were incorrect (40.91%), and 39 were correct (59.09%). Here, it appears that the simpler layout, with fewer points of interest, leads to better accuracy. Again, no evident relation between the query category and the percentage of incorrect instructions can be observed.
TABLE I: Results for Bergamo
Query Incorrect Correct
BER-1 (S) 7 (87.50%) 1 (12.50%)
BER-2 (M) 7 (70.00%) 3 (30.00%)
BER-3 (L) 12 (54.55%) 10 (45.45%)
BER-4 (Spec) 3 (60.00%) 2 (40.00%)
BER-5 (Spec) 4 (50.00%) 4 (50.00%)
Total 33 (62.26%) 20 (37.74%)
TABLE II: Results for Bologna
Query Incorrect Correct
BOL-1 (S) 8 (47.06%) 9 (52.95%)
BOL-2 (M) 4 (28.57%) 10 (71.43%)
BOL-3 (L) 6 (42.86%) 8 (57.14%)
BOL-4 (Spec) 5 (62.50%) 3 (37.50%)
BOL-5 (Spec) 4 (30.77%) 9 (69.23%)
Total 27 (40.91%) 39 (59.09%)
IV-C Orio Center
The map of the Orio Center shopping mall was selected for its more sophisticated and challenging corridor layout. It features more changes in direction, and more corridors that are parallel and perpendicular to each other. Additional complexity arises from smaller retail outlets in the middle of the corridor, thus posing a challenge for the LLM to determine whether these should be considered to the right or left of the user’s path. Unlike the previous tests, all tests were conducted not with IDs to indicate the starting or ending points, but rather with the actual names of the establishments within the map, as this was the information displayed in the image.
TABLE III: User prompt examples for the Bologna Airport.
Scenario ID Query Category User Prompt Query Rationale
Table IV presents the detailed results for Orio Center. Of 67 indications, 32 were incorrect (47.76%), and 35 were correct (52.24%). This indicates an improvement in terms of performance compared to the Bergamo case. The increased intricacy of Orio Center’s layout does not negatively impact the performance of the LLM. Again, the path length does not clearly correlate with performance, as the query of category L (ORIO-3) has the highest percentage of correct indications.
TABLE IV: Results for Orio Center
Query Incorrect Correct
ORIO-1 (S) 5 (71.43%) 2 (28.57%)
ORIO-2 (M) 5 (50.00%) 5 (50.00%)
ORIO-3 (L) 4 (22.22%) 14 (77.78%)
ORIO-4 (Spec) 13 (76.47%) 4 (23.53%)
ORIO-5 (Spec) 5 (33.33%) 10 (66.67%)
Total 32 (47.76%) 35 (52.24%)
IV-D Overall Performance
The overall performance across all locations is summarized in Table V. A total of 186 output indications were evaluated, with 92 incorrect (49.64%), and 94 correct (50.54%). This indicates that, while improvements are certainly required, the performance of the LLM is promising, also considering the complexity of the different scenarios and paths. While the best performance is obtained for the “simplest” map, i.e., Bologna Airport, challenging scenarios like Orio Center do not appear to affect the performance strongly. Instead, for a case like Bergamo, which also has a simple layout with a single corridor, performances are significantly lower. We argue that this is due to the high number of points of interest and more icons, which appear to overwhelm the LLMs. Our results suggest that it is not the layout or the complexity of the path that affects the performance, but rather the amount of information reported on the map. This can be improved by providing different, simplified views of a map to be used as input to the LLM to facilitate the task. Another phenomenon that we observed, which we can relate to the map’s visual appearance, is the types of errors made for the Orio Center case. In some situations, the LLM gave indications of routes that were not physically possible due to the presence of walls, which were not clearly recognizable due to the similar shades of colours on the map. We thus argue that an image with more contrasting colours could lead to better performance. This can be done in a pre-processing phase, together with removing the legend. It should be highlighted that the maps were all coloured, and in future works, we plan to evaluate the performance considering high-contrast greyscale images, which could simplify the task by discarding irrelevant colour information.
Another phenomenon we observed is that, in many cases, such as the one exemplified in Section III, the LLM seems to struggle with adopting the user’s perspective. Instead of providing directions from the point of view of a subject within the corridors, the output appears to be based on the perspective of an external observer looking at the map. These issues could, in principle, be avoided by identifying the user’s viewpoint and position, which can be achieved by asking them to describe what they see to their left, right, and in front of them.
TABLE V: Overall Results
Location Incorrect Correct
Bergamo 33 (62.26%) 20 (37.74%)
Orio Center 32 (47.76%) 35 (52.24%)
Bologna 27 (40.91%) 39 (59.09%)
Total 92 (49.46%) 94 (50.54%)
IV-E Comprehensive Evaluation of Incorrect Answer
In the comprehensive evaluation of the responses generated by the model during testing, it was discovered that 49.46% were incorrect. To understand the nature of these inaccuracies, a detailed analysis was undertaken. It became clear that these incorrect responses were not in accordance with the expected guidelines. However, this analysis also yielded valuable insights, revealing that the guidance provided does not negatively affect the user’s overall direction and progress despite the errors.
An example, taken from the response related to Fig.  3, of insights categorized in this analysis as ’Partially Correct’ includes:
  • “… which will also be on your left.” where the guidance is incorrect, but it is not crucial and does not harm the user’s journey. It is merely a wrong choice of reference point along the path. Errors like this can be classified as less severe than others.
  • “and you will see Clayton” Where there was a mistake in the order of two points of interest along the route. According to the directions, the user should see Dainese first and then Clayton, but in this case, the model reversed their order. This, too, is a minor error that does not compromise the overall direction of the user.
The set of these indications considered partially correct constitutes 16% of the total incorrect responses.
In many cases, as seen in this example, there is a sense that the LLM fails to put itself in the user’s perspective (even when specified). The output does not refer to the point of view of a virtual subject within the corridors but gives directions based on the perspective of an observer looking at the map. These issues can, in principle, be prevented by identifying the position of users, which can be done by asking them to describe what they see on their left/right and in front of them.
V Conclusions
This paper evaluates the performance of ChatGPT in providing human-like indications to users in indoor environments represented through visual maps. We considered three different real-world map layouts with different characteristics, and five types of engineered user prompts of different complexity levels. Our preliminary results are promising, with an average of 50.54% correct indications, and a maximum of 77.78%. Furthermore, we noticed that the complexity of the user query and of the layout of the map do not appear to strongly influence the performance. Instead, these appear to be affected by the visual characteristics of the maps, e.g., the degree of contrast between colours and size and the information overload due to the density of points of interest. In future work, we plan to consolidate and improve these results with more systematic tests, including pre-processed maps, e.g., with increasing contrasts and simplified information. Furthermore, to go towards a more practical application of the envisioned approach, we plan to pre-process natural user queries so that these can be translated into engineered user prompts, similar to the ones used in our tests. These improvements aim to make the system more robust and user-friendly, paving the way for practical applications in real-world indoor navigation systems.

### v4

@@ III-A Evaluation Method
IV Experimental Results
This section presents and analyzes three individual experiments in detail. For each experiment, the complete pipeline is described step by step. Summary tables provide a comprehensive overview of success and failure rates across different queries and map configurations. The section concludes with a detailed discussion of common errors encountered and the improvements made over the previous version of the pipeline.
IV-A Experiment I - Original Image Map
The first experiment was performed with ChatGPT 4o. The complete and detailed analysis of this experiment is described in [10].
  • Pre-processing: Removed the legend from the image, to ensure that the input map is free of noise that can affect the LLM performance.
  • System prompt: defined as ’SP#1’, was designed based on the characteristics of the input image. Before data collection, several improvements were made to the prompt to enhance the model’s overall performance. After various tests, it became evident that the LLM’s output exhibited recurring errors, e.g., evaluating the proximity between points based on the numerical difference between their IDs on the map. These prompt refinements were introduced specifically to address those issues and to better guide the model’s reasoning in relation to the visual context.
In Table II, a total of 186 output indications were evaluated, with 92 incorrect (49.64%), and 94 correct (50.54%). The highest score of correct indications is for Bologna Airport map (59.95%) and the lowest result is for Bergamo Airport map (37.74%). These results suggest that clearer and less cluttered visual representations lead to better model performance. Notably, the Bologna map contains fewer facilities and additional information, which likely contributed to the improved navigation accuracy.
TABLE II: Results with Image without legend and System prompt SP#1 with ChatGPT-4o
Query BERGAMO BOLOGNA ORIO
Incorrect Correct Incorrect Correct Incorrect Correct
Query-1 (S) 7 (87.50%) 1 (12.50%) 8 (47.06%) 9 (52.95%) 5 (71.43%) 2 (28.57%)
Query-2 (M) 7 (70.00%) 3 (30.00%) 4 (28.57%) 10 (71.43%) 5 (50.00%) 5 (50.00%)
Query-3 (L) 12 (54.55%) 10 (45.45%) 6 (42.86%) 8 (57.14%) 4 (22.22%) 14 (77.78%)
Query-4 (Spec) 3 (60.00%) 2 (40.00%) 5 (62.50%) 3 (37.50%) 13 (76.47%) 4 (23.53%)
Query-5 (Spec) 4 (50.00%) 4 (50.00%) 4 (30.77%) 9 (69.23%) 5 (33.33%) 10 (66.67%)
Total 33 (62.26%) 20 (37.74%) 27 (40.91%) 39 (59.09%) 32 (47.76%) 35 (52.24%)
Furthermore, we noticed that the complexity of the user query and the layout of the map do not appear to strongly influence the performance. Instead, these appear to be affected by the visual characteristics of the maps, e.g., the degree of contrast between colours and size and the information overload due to the density of POIs.
IV-B Experiment II - Graph Style Map
This experiment aims to enhance the performance of the LLM, particularly in identifying the correct path and accurately recognizing POIs and other facilities based on the user’s perspective. It specifically addresses the most frequent errors observed in the previous experiments.
  • Pre-processing: the original images from Experiment I were manually replaced with a simplified graphical representation consisting solely of corridors and POIs as in the Figure 3. Corridors are depicted as black lines, while POIs are represented by green dots labeled with their corresponding ID or name. This was carried out to increase the readability of the image by the model. The use of numbers, characters, or words to identify POIs in the new graph-style map reflects the conventions of the original maps. Legends using alphanumeric references were preserved as such, while named locations remained textual. Icons were converted into verbal descriptions
  • System prompt: SP#1, same as the previous Experiment I).
Figure 3: Graph-style image of Orio Center mall after Pre-processing phase.
TABLE III: Results with graph-style images and System prompt SP#1 with ChatGPT-4o
Query BERGAMO BOLOGNA ORIO
Incorrect Correct Incorrect Correct Incorrect Correct
Query-1 (S) 0 (0.00%) 7 (100.00%) 4 (40.00%) 6 (60.00%) 4 (57.14%) 3 (42.86%)
Query-2 (M) 1 (16.67%) 5 (83.33%) 2 (14.29%) 12 (85.71%) 7 (43.75%) 9 (56.25%)
Query-3 (L) 5 (33.33%) 10 (66.67%) 4 (23.53%) 13 (76.47%) 9 (40.91%) 13 (59.09%)
Query-4 (Spec) 1 (20.00%) 4 (80.00%) 1 (12.50%) 7 (87.50%) 1 (10.00%) 9 (90.00%)
Query-5 (Spec) 6 (27.27%) 16 (72.73%) 1 (5.56%) 17 (94.44%) 6 (35.29%) 11 (64.71%)
Total 13 (23.64%) 42 (76.36%) 12 (17.91%) 55 (82.09%) 27 (37.50%) 45 (62.50%)
In Table III, a total of 194 output indications were evaluated, with 52 incorrect (26.80%), and 142 correct (73.20%). The highest score of correct indications set is for Bologna Airport map (82.09%) and the lowest result is for Orio-Center map (62.50%).
There is a clear improvement, from which it can be deduced that stylized, graph-like maps perform better, thanks to the more essential (and relevant) information included. The major improvement observed in these tests is that the logical path through the corridors is followed correctly. All POIs along the route are correctly identified and presented. The most frequent errors still concern orientation: the model often fails to identify the initial direction to take, or whether a POI is located to the right or left. This could still be a problem of map representation, or it might be an intrinsic logical limitation of the LLM’s internal reasoning, preventing it from correctly interpreting directionality.
IV-C Experiment III - Model Upgrade
This pipeline was initially tested using the same language model employed in the previous two experiments. However, this model did not provide strong evidence of a significant improvement over Experiment II. Therefore, we opted for the most powerful model of OpenAI, known as ChatGPT-o3.
  • Pre-processing: Same images as Experiment II, with the graph representation based on corridors and POI.
  • System prompt: defined as ’SP#2’. The difference between this and SP#1 is only the proper and clear description of the map given by input.
TABLE IV: Results with graph-style images and system prompt SP#2 using ChatGPT-o3 (with reasoning)
Query BERGAMO BOLOGNA ORIO
Incorrect Correct Incorrect Correct Incorrect Correct
Query-1 (S) 1 (10.00%) 9 (90.00%) 2 (18.18%) 9 (81.82%) 1 (11.11%) 8 (88.89%)
Query-2 (M) 1 (5.88%) 16 (94.12%) 2 (6.67%) 2 (93.33%) 2 (8.70%) 21 (91.30%)
Query-3 (L) 1 (2.86%) 34 (97.14%) 8 (28.57%) 20 (71.43%) 7 (18.42%) 31 (81.58%)
Query-4 (Spec) 2 (28.57%) 5 (71.43%) 2 (13.33%) 13 (86.67%) 1 (5.26%) 18 (94.74%)
Query-5 (Spec) 4 (14.81%) 23 (85.19%) 5 (15.15%) 28 (84.85%) 5 (19.23%) 21 (80.77%)
Total 9 (9.38%) 87 (90.63%) 19 (16.24%) 98 (83.76%) 16 (13.91%) 99 (86.09%)
In Table IV, a total of 328 output indications were evaluated, with 44 incorrect (13.41%), and 284 correct (86.59%). The highest average score of correct indications set is for Bergamo Airport map (90.63%) and the lowest result is with Bologna Airport map (83.76%). With a peak of performance up to 97.14% in Ber-3 query.
A relevant increase in the number of correct navigational indications was observed, particularly with a higher percentage of correctly identified left and right turns. The model demonstrated an improved sense of orientation throughout the path, accurately signaling directions all along the path.
Our findings demonstrate that the model achieves high levels of accuracy and reasoning performance, significantly outperforming the first two experiments.
V Discussion and conclusion
TABLE V: Overall average results over tests
Experiment Incorrect Correct
Experiment I 49.46% 50.54%
Experiment II 26.80% 73.20%
Experiment III 13.41% 86.59%
This paper evaluates the performance of ChatGPT in providing human-like indications to users in indoor environments represented through visual maps. We considered three different real-world map layouts with different characteristics, five types of engineered user prompts of different complexity levels, and three experimental settings, considering two different LLMs and pipeline strategies. In the three experiments, a total of 708 output indications were evaluated. The overall performance across all different experiments is summarized in Table V. The results indicate that, with ChatGPT4-o3 and a stylized representation of the indoor maps (Experiment III), we can achieve 86% correct indications, which is a highly promising result.
Given the challenges of indoor navigation, this research has several key implications for AI-driven navigation, human-computer interaction, and assistive technologies. Users often ask questions such as “What can I find on this floor?” or “Where is the nearest restroom?” highlighting the need for adaptable navigation solutions that can dynamically respond to different environments and user preferences. A detailed and accurate description of the surrounding environment can enhance spatial awareness and improve the overall navigation experience. The use of LLMs can substantially reduce the costs of an indoor navigation system, as it relies solely on software. Furthermore, the LLM requires only a pre-processed, graph-like map of the target building, which speeds-up the deployment in diverse environments, even in the presence of layout updates (e.g., due to work-in-progress, or expansion of the building). In our envisioned system, users can use their mobiles to interact with an LLM-based assistant, which can guide them to their desired destination efficiently, especially if the LLM can also leverage wi-fi fingerprints and provide instructions considering the current position of the user. The use of wi-fi can also help to correct the errors introduced by the LLMs (e.g., confusion between left and right), thereby improving the quality of the indications.
A critical limitation of the best solution identified is the average response time—approximately five minutes per query—posing a substantial barrier to real-time usage. Future research directions include exploring more optimized inference strategies or lightweight alternatives to reduce latency without compromising the quality of reasoning. Another challenge to address is the automatic translation of the user query into an engineered prompt, which can, in principle, be achieved again through the mediation of the LLM.
Besides addressing the current limitations, our future work will be oriented to evolve the idea to support users with different accessibility problems, also through the integration with localization systems.
Future iterations could integrate real-time obstacle detection via smartphone sensors or connected assistive devices (e.g., smart canes) to supplement LLM-generated instructions. For individuals with motor disabilities, the system can prioritize accessible routes and adapt instructions based on mobility aids.
Additionally, voice-controlled navigation and hands-free querying can reduce physical interaction demands.
In conclusion, our paper is the first step to make indoor navigation not only similar to outdoor navigation systems in terms of user experience, but also more inclusive towards diverse accessibility needs while reducing hardware dependencies.